# Christian Peukert [^1] Corresponding author. Florian Abeillon Jérémie Haese Franziska Kaiser Alexander Staub

### v1

@@ 7 Conclusion
GenAI models benefit from the abundance of available training data on the internet. However, the economics of the supply of data have not been discussed much. In this research, we focus on how strategic behavior of human creators affects the flow and, therefore, the future stock of AI training data.
We study the empirical setting of creators who openly contribute their work to Unsplash, a stock photography platform that provides crowd-sourced images under a permissive license. In particular, we evaluate the effects of the release of a training dataset for commercial AI applications. We find that affected users are more likely to leave the platform. Those that stay, reduce their uploads substantially. As a result, the size of the flow of data changes, as does the content. Our results imply that the release of a free training dataset for commercial AI applications, covering roughly 5% of the platform’s contributors, overall reduced the flow of data by 2%, decreased the variety (deviation from the average concept in the stock) in the dataset by 4% and increased novelty (repetition of very similar concepts in the stock) by 2.5% over the course of three years.
Our results have important implications for the debate on copyright and AI. Even in a setting where copyright does not directly factor in – Unsplash caters to a special set of creators that choose an open license –, we see a sizable change in individuals’ behavior when their work is made available as an AI training dataset. Our results suggest that a change in the flow of data can also translate into changes in aspects of dataset quality, such as variety and novelty, because of changes in the user composition. Consequently, this has implications for the quality of AI training datasets and possibly AI output (Hadsell et al., 2020; He et al., 2011; Quinn and Gutt, 2023; Whang et al., 2023).
In conclusion, our results raise the question of whether exemptions in copyright law for AI (e.g. through fair use provisions that do not compensate rightsholders) could lead society to a situation of lower quality AI (smaller and less varied datasets) and less interesting (because less varied and more repetitive) human creations. Hence, while AI developers may initially prefer laissez-faire copyright (e.g., fair use or text-and-data-mining exceptions), the upstream consequences on the flow of data, both in terms of quantity and quality, may have negative consequences for the continued quality of existing AI applications and raise difficulties in the development of new technologies.
More research is needed to guide policymakers to carefully weigh the trade-offs between regulation of the supply of data, e.g. through privacy policy, competition policy, or copyright law, and AI innovation.
@@
Note: We estimate Poisson pseudo-maximum likelihood regressions models. The dependent variable is the number of uploads in a given month. Treated indicates whether a user was part of the LITE dataset. Above99thPct, Above95thPct, and Above90thPct indicate whether a user’s uploads were above the 99/95/90th percentile. The samples in columns (3) to (5) do not include users above the 99/95/90th percentile of the pre-treatment uploads. The number of observations is smaller than in Tables 3 and 5 because singleton observations are dropped due to fixed effects. Standard errors clustered at the user level in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$ $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$

### v2

@@ 6 Conclusion
We investigate the behavioral reactions of creators when their works are made available for the training of commercial AI applications and how this affects the continued supply of content. Our empirical context provides us with a natural experiment on the photography platform Unsplash we exploit to derive empirical insights. Our results imply that the release of an AI training dataset for commercial use, covering roughly 5% of the platform’s contributors, reduced overall uploads (and, therefore, the flow of AI training data) by 2%. It also decreased the variety (deviation from the average image in the stock) of the average upload by 5%, and the novelty of the average upload (repetition of very similar images in the stock) by 14-28%.
Our findings echo prior work documenting adverse reactions from content creators when AI is introduced into their domains (Huang et al., 2023; Lin, 2024; Quinn and Gutt, 2025) and correspond to the broader reactions of rights holders, including writers, musicians, and visual artists.
Prior literature has focused on the demand-side implications of AI (Felten et al., 2023; Yiu et al., 2024; Brynjolfsson et al., 2025; Demirci et al., 2025), whereas studying the upstream supply of data forms the contribution of our research. We present empirical evidence for theoretical arguments on AI-induced market changes (Wang et al., 2024; Yang and Zhang, 2024; Gans, 2025), and provide the microfoundations – causal evidence on the behavioral responses of data contributors – for the broader trend in the decreasing availability of open data on the web (Longpre et al., 2023). Our research complements previous empirical studies on how changes in the supply of human-generated data may impact the quality of AI training datasets (Jia and Nagaraj, 2025). Understanding these dynamics is crucial as AI development is moving forward in an environment where high-quality, diverse, and legally accessible data may become increasingly scarce.
The computer science literature suggests that synthetic data – AI-generated rather than human-created – could reduce the reliance on human-generated content (Ho et al., 2024). However, studies warn that excessive use of synthetic data in training can degrade model performance, leading to what is known as “model collapse” (Valavi et al., 2022b; Shumailov et al., 2024). Hence, technical solutions seem difficult, which implies that setting the right incentives for human creators to continue contributing to AI training datasets forms a first-order problem for managers and policymakers. The solutions need to include compensation mechanisms and data usage frameworks that align the incentives of creators, content platforms, AI developers, and the users of AI models. Concerning the aspects mentioned above, the current policy appears misguided. For example, the existing regulatory framework in Europe requires high-quality training datasets that are compliant with copyright law, which allows rights holders to opt out of certain uses of their works. [^35] See EU AI Act, Article 10(3), https://eur-lex.europa.eu/legal-content/EN/TXT/?uri=OJ:L_202401689#art_10 and EU Copyright Directive, Article 4, https://eur-lex.europa.eu/legal-content/EN/TXT/?uri=CELEX:32019L0790#art_4. Our findings highlight that such a regime can introduce bias in training datasets and lower the quality of AI model outputs.
Our paper provides a first step to understanding the supply side of AI training data, but further research will be required moving forward. From a technical standpoint, we need direct evidence on how a decrease in the flow of data – specifically in its quantity, variety, and novelty – affects AI model performance. From a practical standpoint, additionnal empirical evidence on how the compensation of creators influences the quantity and quality of training data to design effective compensation mechanisms is of high importance.
@@
[…]
2import os
3import time
4import requests
5import pandas as pd
6from tqdm import tqdm
7from concurrent.futures import ThreadPoolExecutor, as_completed
8
9file_path = "unsplash-research-dataset-lite-1.0.1/photos.tsv000"
[…]
11output_dir = "unsplash-research-dataset-lite-1.0.1/images"
12os.makedirs(output_dir, exist_ok=True)
13
14# Function to download a single image
15def download_image(row):
16 image_url = row.get("photo_image_url")
17 photo_id = row.get("photo_id")
18
19 if pd.notna(image_url) and pd.notna(photo_id):
20 try:
21 response = requests.get(image_url, stream=True, timeout=10)
22 if response.status_code == 200:
23 image_path = os.path.join(output_dir, f"{photo_id}.jpg")
24 with open(image_path, "wb") as f:
25 for chunk in response.iter_content(1024):
26 f.write(chunk)
27 return True
28 except Exception:
29 pass
30 return False
31
32max_threads = 10
33print("Starting image downloads in parallel...")
34
35start_time = time.time()
36
37with ThreadPoolExecutor(max_workers=max_threads) as executor:
38 futures = [executor.submit(download_image, row) for _, row in df.iterrows()]
39 for _ in tqdm(as_completed(futures), total=len(futures)):
40 pass
41
42elapsed_time = time.time() - start_time
43print(f"Downloaded {len(df)} images in {elapsed_time:.2f} seconds."
Note: Python script to download all images released in the LITE version 1 dataset. With this script below, the authors were able to download 24,989 images of the first version of the LITE dataset (81.04GB) in 21 minutes on a standard MacBook Pro (11 images were no longer available).
[…]
2SELECT DISTINCT photos.id FROM photos
3JOIN tags ON tags.photo_id = photos.id
4WHERE tags.keyword = nature -- The image is tagged with "nature"
6 AND photos.curated = TRUE -- image has been curated by the editorial team
7 AND photos.status = 1 -- image is still available on the platform
8LIMIT 25000
Note: This query was sent to the authors by the head of data science at Unsplash.
Table A.1: Results: Likelihood of remaining on the platform (interacted)
Images
(1) (2) (3)
Treated 0.0109∗∗∗ 0.0226∗∗∗ 0.0203∗∗∗
(0.00155) (0.00364) (0.00401)
Treated $\times$ Pro-Gear -0.0051 -0.0067
(0.00329) (0.00345)
Treated $\times$ AccountAge -0.0003∗∗∗ -0.0002∗∗∗
(0.00008) (0.00008)
Treated $\times$ Multiple -0.0066∗∗∗ -0.0071∗∗∗
(0.00243) (0.00241)
Treated $\times$ ForHire 0.0091∗∗∗
Mean Control 0.9075 0.9075 0.9075
Image-Age FE Yes Yes Yes
User FE Yes Yes Yes
Observations 704,848 704,848 648,653
Note: We estimate a linear probability model. The dependent variable is an indicator for whether an image remains on Unsplash as of May 2023. The sample in column (3) is smaller because data on “ForHire” is not available for all users. The variable Treated indicates whether a user was part of the LITE dataset. We compare survival rates over 2.5 years (from Aug 2020 to May 2023). We include fixed effects for the image’s age (in months) in all specifications. Standard errors clustered at the user level in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
Table A.2: Results: Changes in uploads (Weighted)
Uploads I(Uploads$>$0) Log(1+Uploads)
(1) (2) (3)
Post $\times$ Treated -1.0573∗∗∗ -0.0685∗∗∗ -0.1599∗∗∗
(0.08746) (0.00351) (0.00715)
Mean Treated Before 2.9038 0.2635 0.4892
Month FE Yes Yes Yes
User FE Yes Yes Yes
Observations 612,662 612,662 612,662
@@
Month FE Yes Yes Yes Yes Yes Yes
User FE Yes Yes Yes Yes Yes Yes
Table A.4: Results: Changes in uploads, alternative control group definitions (weighted)
Not Nature&Curated Not in LITE
(1) (2) (3) (4) (5) (6)
Uploads I(Upl.$>$0) Log(1+Upl.) Uploads I(Upl.$>$0) Log(1+Upl.)
Post $\times$ Treated -1.3176∗∗∗ -0.0849∗∗∗ -0.2037∗∗∗ -1.3930∗∗∗ -0.1033∗∗∗ -0.2287∗∗∗
(0.07575) (0.00236) (0.00525) (0.07567) (0.00235) (0.00524)
Mean Treated Before 0.4892 0.2635 0.4892 0.4892 0.2635 0.4892
@@
Note: We estimate Poisson pseudo-maximum likelihood regressions models. The dependent variable is the number of uploads in a given month. Treated indicates whether a user was part of the LITE dataset. Above99thPct, Above95thPct, and Above90thPct indicate whether a user’s uploads were above the 99/95/90th percentile. The samples in columns (3) to (5) do not include users above the 99/95/90th percentile of the pre-treatment uploads. The number of observations is smaller than in Tables 3 and A.6 because singleton observations are dropped due to fixed effects. Standard errors clustered at the user level in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$ $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
Table A.6: Results: Changes in uploads, by user activity
(1) (2) (3) (4)
All w/o 99th w/o 95th w/o 90th
Post $\times$ Treated -0.4529∗∗∗ -0.9872∗∗∗ -0.6847∗∗∗ -0.4571∗∗∗
(0.02672) (0.05445) (0.03513) (0.02699)
Post $\times$ Treated $\times$ Above99thPct -17.4338
Post $\times$ Treated $\times$ Above95thPct -0.9941
Post $\times$ Treated $\times$ Above90thPct -1.3862
Mean Treated Before 2.9038 2.2160 1.5488 1.1435
Month FE Yes Yes Yes Yes
User FE Yes Yes Yes Yes
Observations 612,662 606,490 581,988 551,076